OpenStudy (twizttiez):

1. Grandma Edith uses her big punch bowl at family gatherings. The bowl is half of a sphere, with a diameter of 30 cm. Her ladle is also half of a sphere, with a diameter of 8 cm. How many ladles full of punch can Grandma Edith scoop from the punch bowl? Use 3.14 to approximate pi and express your answer as a whole number. Show your work. Can someone walk me through this?

OpenStudy (twizttiez):

@campbell_st

OpenStudy (campbell_st):

ok.... so the volume of a sphere is \[V = \frac{4}{3} \pi r^3\] you need to find the radius of the bowl... then find the volume of the bowl, as a sphere... a hemisphere is 1/2 a sphere, so have the answer for the volume of the bowl. same formula and process for the ladle.. 1. find the radius... 2. find the volume of the sphere 3. halve the answer for the volume of the ladle. last part... number of ladles = volume of bowl / volume of ladle hope it helps

OpenStudy (twizttiez):

Thanks so much

OpenStudy (campbell_st):

glad to help

OpenStudy (twizttiez):

Wait okay so would I do 4/3*3.14*30^3=113040 113040/2=56520 to get the volume of the bowl?

OpenStudy (twizttiez):

That's what I did and that's what I got

OpenStudy (twizttiez):

????

OpenStudy (twizttiez):

So I did all of that and my final answer came to 52.734475 and I rounded it to a whole number and got 52 obviously am I correct? ^-^

OpenStudy (campbell_st):

no, the question gives diameters... you need to halve the diameter to get the radius... diameter 30 cm then radius r = 15 diameter 8 cm then radius r = 4 cm so but you'll need to recalculate the volumes.

OpenStudy (twizttiez):

Ohhhhh

OpenStudy (twizttiez):

Okay :"(

OpenStudy (campbell_st):

sorry about that... halve each volume to get the hemisphere...

OpenStudy (twizttiez):

Okay I understand so do what I did just use different numbers? XD

OpenStudy (twizttiez):

So this time use 4/3*3.14*15^2=942 942/2=471?

OpenStudy (campbell_st):

ok... here is what you need Punch bowl.... and I'll halve the volume in the process \[V =\frac{1}{2} \times( \frac{4}{3} \times 3.14 \times 15^3)\] ladle... all in 1 volume \[V = \frac{1}{2} \times ( \frac{4}{3} \times 3.14 \times 4^3)\] hope it helps..

OpenStudy (twizttiez):

OHHHHHHHHH

OpenStudy (twizttiez):

Wait that's what I did just longer XD

OpenStudy (twizttiez):

But I see it's a more simplified way thanks!

OpenStudy (campbell_st):

you just had the power wrong... things were correct otherwise.. so calculate the volumes then its bowl divided by ladle... rough guess is something around 50...

OpenStudy (campbell_st):

yes, I tried to step you though it... rather the give you the whole thing, both methods are correct.

OpenStudy (twizttiez):

Yup yup

OpenStudy (twizttiez):

I have one more question could you help me out?